OpenStudy (anonymous):

(4/x)+(12/x-5)=-11 solve its has two solutions can someone show me how to solve this step by step

OpenStudy (jit4won):

yes

OpenStudy (jit4won):

multiply each term by x

OpenStudy (anonymous):

i see how yougot the denominator but that numerator i dont get it

OpenStudy (jit4won):

wait

OpenStudy (jit4won):

There is an easy way

OpenStudy (jit4won):

Look If you have equation like a/b + c/d = e just cross multiply the denominators That is (a x d + c x b)/ bx d

OpenStudy (jit4won):

here you can write (4 ( 12 -x ) + 12 X )/ X ( 12-X)= - 11

OpenStudy (jit4won):

WHICH GIVES YOU (48 - 4X + 12 X )/ 12X- X^2 = -11

OpenStudy (jit4won):

48 + 8X / 12X- X^2 = -11

OpenStudy (jit4won):

48 + 8X = - 11( 12X - X^2)

OpenStudy (jit4won):

48 + 8X = -132X + 11 X^2 11X^2 -132X -8X -48 = 0 11X^2 -140X -48 = 0

OpenStudy (jit4won):

X = - b ± √(b2 - 4ac) / 2a

OpenStudy (jit4won):

a = 11 b= -140 c= - 48